OpenStudy (anonymous):

Find dy / dx by implicit differentiation: 5 + 3 x = sin (x y^7)

OpenStudy (anonymous):

\(\large 5 + 3 x = sin (x y^7) \) \(\large [5 + 3 x]' = [sin (x y^7)]' \) \(\large [5]' + [3 x]' = cos (x y^7) \cdot [xy^7]' \) \(\large 0+3 = cos (x y^7) \cdot [xy^7]' \)

OpenStudy (anonymous):

@monroe17 Now your next step is applying product rule: ( xy^7) ' = ...

OpenStudy (anonymous):

sorry.... afk for a bit... yes, can you do the product rule on [xy^7]' ???